OpenStudy (anonymous):

Evaluate the expression for x = –3.

OpenStudy (hba):

Which expression ?

OpenStudy (abb0t):

plug in 3 in place of x. \[e-3pression\]

OpenStudy (hba):

Its -3 @abb0t

OpenStudy (anonymous):

X squared + 4x + 3

OpenStudy (hba):

x^2+4x+3 Just plug in -3 in place of x :) Try solving and tell me what you get @Tekoa

OpenStudy (anonymous):

\[(-3)^{2}+4(-3)+3\]is what @Tekoa

OpenStudy (anonymous):

@hba i got 18

OpenStudy (anonymous):

@Tekoa What is\[(-3)^{2}\]?

OpenStudy (anonymous):

@aylin i dont know :(

OpenStudy (hba):

I guess you are making sign problem. (+)(-)=+ (-)(-)=+ (-3)^2=(-3)(-3)=

OpenStudy (anonymous):

\[(-3)^{2}\]is the same thing as\[-3 \times -3\]

OpenStudy (hba):

(+)(-)=-*

OpenStudy (anonymous):

oh lord !!!!

OpenStudy (anonymous):

@Aylin oh

OpenStudy (anonymous):

So with what hba said,\[-3 \times -3 =9\]Also\[4 \times -3 = -12\]So\[(-3)^{2}+4(-3)+3=-3 \times -3 + 4 \times -3 +3=9-12+3=?\]So what does the last one equal?

OpenStudy (anonymous):

@Aylin i dont know , im not good at math

OpenStudy (anonymous):

Which of the following verbal expressions represents the algebraic expression 2x – 8?

OpenStudy (anonymous):

eight less than twice a number B. the difference between eight and twice a number C. twice eight less than a number D. twice a number and eigh